# Disentangling Visual Transformers: Patch-level Interpretability for Image Classification

### v1

@@ 1 Introduction
To encourage future research, we will release our code and pretrained weights upon publication.
@@ 2 Related Work
Concerning visual transformers, several recent works have attempted to make transformers more interpretable by modifying their architecture. Some papers try to push the boundaries to include transformer-based heads [25, 48, 51, 31] for prototype interpretability. However, they only include a single self-attention mechanism on top of a CNN backbone. ProtoPFormer [68] suggests including a ViT backbone in the prototype setup. To do this, ProtoPFormer includes a prototype layer on top of both the classification and image tokens. However, this architecture does not guarantee that the tokens contain purely local information, especially since computing a classification on top of image tokens increases the receptive fields of the tokens [50]. While innovative, this architecture does not ensure that the tokens contain purely local information due to the increased receptive fields of tokens when combining classification and image tokens [50]. B-cos networks V2 [10] use the same rational as the original B-cos [7] approach, but they extend it to transformers. In a few words, B-cos networks summarize their inner workings as a single linear function, creating the attribution map by simply multiplying the input and the summarized network. However, this extension still relies on convolutional layers as the initial feature extraction stage, limiting its interpretability scope. Finally, A-ViT [69] was originally designed for faster inference by removing tokens at certain layers. Despite this, this mechanism serves as an interpretable system, as the most important tokens are retained until the last layer. This approach, however, focuses on efficiency rather than interpretability as its primary goal.
@@ 3.1 Preliminaries: Transformers, ViTs and Notations
The transformer architecture is built upon the Scaled Dot-Product Attention operation [63], commonly referred to as the attention. Given a query token sequence ${x}^{q}\in\mathbb{R}^{L_{q}\times d_{model}}$ and a target sequence (or key-value sequence) ${x}^{t}\in\mathbb{R}^{L_{t}\times d_{model}}$, where $L_{q}$ and $L_{t}$ are their respective sequence lengths and $d_{model}$ is the token dimension, the attention mechanism is computed as follows:
@@
$MHA(x^{q},x^{t})=\underbrace{[A^{1}(x^{q},x^{t});...;A^{h}(x^{q},x^{t})]}_{\texttt{Concatenate $h$ times}}W_{o}+b_{o},$ (2)
@@
In computer vision, to incorporate image data into this sequence-based formulation, the ViT first partitions the input image into $N^{2}$ equal-sized patches and linearly projects them to create the patch token sequenceFor the rest of the paper, we will use the terms token and patch interchangeably, referring to the image patch tokens.. Additionally, following standard practice, a learnable classification token CLS is prepended to the patch sequence. Furthermore, each patch token is summed with a positional embedding to encode its spatial location within the image. For the remainder of the paper, the sequence $x\in\mathbb{R}^{(N^{2}+1)\times d_{model}}$ denotes the concatenation of the patch tokens and the CLS token, where $x[0]$ corresponds to the CLS token.
@@ 3.2 Multi-Head Attention and Patch Mixing in Transformers
$A(x^{q},x^{t})=\sum_{v\in x^{t}}a(v,x^{q},x^{t})(v\,W_{V}+b_{v}),$ (4)
where $a(v,x^{q},x^{t})$ is the attention of a single token $v\in x_{t}$. Here, Eq. 4 shows that we can decompose the attention mechanism into separately processed patches - each patch $v$ in $x^{t}$ adds $a(v,x^{q},x^{t})(v\,W_{V}+b_{v})$. Accordingly, if $x^{t}$ contains purely local information, the output of the attention is a sum of local data.
@@
$\begin{split}MHA(x^{q},x^{t})&=b_{o}+\sum_{v\in x^{t}}v^{\prime}(v)\\ \text{where}\quad v^{\prime}(v)&=\sum_{i=1}^{h}a(v,x^{q},x^{t})(v\,W_{v}^{i}+b_{v}^{i})W_{o}^{i}.\end{split}$ (5)
@@ 3.3 Untangling Visual Transformers
One advantage of this architecture is that we can easily compute saliency maps, as shown in Fig. 2. The double sum in Eq. 8 can be decomposed as a tensor $\mathrm{R}^{L\times N^{2}\times d_{model}}$, where the final image representation is the sum over the first and second dimensions, i.e. the layer and token dimension, respectively. Thus, and similarly to CAM [73], to compute the regions of interest used by the model for an input image, we simply run the linear classifier on each patch to get the map. This rationale is similar to the LRP [6] method in the sense that the sum of value in the saliency is equal to the output logit for that specific class.
@@ 3.4 Token Pooling
Token pooling [41], which involves downsampling the number of tokens as one progresses through the layers, is commonly used to improve the computational efficiency of standard transformers. This pooling technique effectively addresses the issue of representation power (as empirically demonstrated in § 4.7) by expanding the receptive field of the image tokens in the deeper layers. We choose to adopt this approach due to its significant advantages.
@@ 4.2 Implementation Details
To train HiT on ImageNet [16], we used the official DeiT3 codebase [60] and followed a similar setup to their method. Model-wise, HiT uses the same configuration as a ViT except that we removed the last MLP block as it is not used. Regarding the pooling layers, we added two. For HiT-B and HiT-S, we trained our models for 600 epochs using the AdamW optimizer [39] with a learning rate of $8\times 10^{-4}$, a weight decay of 0.05, a batch size of 4096, 20 warm-up epochs, a cosine annealing scheduler [38], and ThreeAugment [60] data augmentation. Unlike the DeiT3 training regime, we did not use the binary cross entropy loss or any LayerDrop [21] regularization, but the traditional cross entropy with a smoothing of 0.1 and an attention dropout of 0.2. To fine-tune HiT in the other datasets, we trained our models similarly to the ImageNet’s configuration, but instead we set the batch size to 512, the number of epochs to 300 and the learning rate to $5\times 10^{-5}$ for Standford Dogs and Oxford Pets, $4\times 10^{-4}$ for Standford Cars and FGVC-Aircraft, and $1\times 10^{-4}$ for CUB 2011.
@@ 5 Conclusions
Appendix A Datasets
@@
Appendix B Insertion-Deletion Curves
@@
Appendix C Evaluating HiT without Pooling Layers
We conducted experiments similar to those in the main manuscript to analyze the positive and negative impact of removing pooling layers. As demonstrated in § 4.7 of the paper, pooling layers are essential for improving top-1 accuracy performance. However, their inclusion increases the size of the explanations. First, we explore this phenomenon quantitatively in sections C.1 and § C.2. Later, we will explore the qualitative differences in § C.3.
C.1 Interpretability Trade-off
@@
C.2 Layer-wise Contributions

### v2

@@ 1 Introduction
Our code and pretrained models are found on GitHub
@@ 2 Related Work
Concerning visual transformers, several recent works have attempted to make transformers more interpretable by modifying their architecture. Some papers try to push the boundaries to include transformer-based heads [25, 49, 52, 31] for prototype interpretability. However, they only include a single self-attention mechanism on top of a CNN backbone. ProtoPFormer [69] suggests including a ViT backbone in the prototype setup. To do this, ProtoPFormer includes a prototype layer on top of both the classification and image tokens. While innovative, this architecture does not ensure that the tokens contain purely local information due to the increased receptive fields of tokens when combining classification and image tokens [51]. B-cos networks V2 [10] use the same rational as the original B-cos [7] approach, but they extend it to transformers. In a few words, B-cos networks summarize their inner workings as a single linear function, creating the attribution map by simply multiplying the input and the summarized network. However, this extension still relies on convolutional layers as the initial feature extraction stage, limiting its interpretability scope. Finally, A-ViT [70] was originally designed for faster inference by removing tokens at certain layers. Despite this, this mechanism serves as an interpretable system, as the most important tokens are retained until the last layer. This approach, however, focuses on efficiency rather than interpretability as its primary goal.
@@ 3.1 Preliminaries: Transformers, ViTs and Notations
The transformer architecture is built upon the Scaled Dot-Product Attention operation [64], commonly referred to as attention. Given a query token sequence ${x}^{q}\in\mathbb{R}^{L_{q}\times d_{model}}$ and a target sequence (or key-value sequence) ${x}^{t}\in\mathbb{R}^{L_{t}\times d_{model}}$, where $L_{q}$ and $L_{t}$ are their respective sequence lengths and $d_{model}$ is the token dimension, the attention mechanism is computed as follows:
@@
$\small MHA(x^{q},x^{t})=\underbrace{[A^{1}(x^{q},x^{t});...;A^{h}(x^{q},x^{t})]}_{\texttt{Concatenate $h$ times}}W_{o}+b_{o},$ (2)
@@
In computer vision, to incorporate image data into this sequence-based formulation, the ViT first partitions the input image into $N^{2}$ equal-sized patches and linearly projects them to create the patch token sequenceFor the rest of the paper, we will use the terms token and patch interchangeably, referring to the image patch tokens.. Additionally, following standard practice, a learnable classification token CLS is prepended to the patch sequence. Finally, each patch token is summed with a positional embedding to encode its spatial location within the image. For the remainder of the paper, the sequence $x\in\mathbb{R}^{(N^{2}+1)\times d_{model}}$ denotes the concatenation of the patch tokens and the CLS token, where $x[0]$ corresponds to the CLS token.
@@ 3.2 Multi-Head Attention and Patch Mixing in Transformers
$A(x^{q},x^{t})=\sum_{v\in x^{t}}s(v,x^{q},x^{t})(v\,W_{V}+b_{v}),$ (4)
where $s(v,x^{q},x^{t})\in\mathbb{R}$ is the contribution factor of $v\in x^{t}$, computed via the softmax operation in Eq. 1. Here, Eq. 4 shows that we can decompose the attention mechanism into separately processed patches - each patch $v$ in $x^{t}$ adds $s(v,x^{q},x^{t})(v\,W_{V}+b_{v})$. Accordingly, if $x^{t}$ contains purely local information, the output of the attention is a weighted sum of local data. Here, we define whether a token contains local information as the product of processing a single patch without the influence of other datapoints.
@@
$\begin{split}MHA(x^{q},x^{t})&=b_{o}+\sum_{v\in x^{t}}v^{\prime}(v)\\ \text{where}\quad v^{\prime}(v)&=\sum_{i=1}^{h}s(v,x^{q},x^{t})(v\,W_{v}^{i}+b_{v}^{i})W_{o}^{i}.\end{split}$ (5)
@@ 3.3 Untangling Visual Transformers
One advantage of this architecture is that we can easily compute saliency maps, as shown in Fig. 2. The double sum in Eq. 8 can be decomposed as a tensor $\mathbb{R}^{L\times N^{2}\times d_{model}}$, where the final image representation is the sum over the first and second dimensions, i.e. the layer and token dimension, respectively. Thus, and similarly to CAM [74], to compute the regions of interest used by the model for an input image, we simply run the linear classifier on each patch to get the map. This rationale is similar to the LRP [6] method in the sense that the sum of value in the saliency is equal to the output logit for that specific class.
@@ 3.4 Token Pooling
Token pooling [41], which involves downsampling the number of tokens as one progresses through the layers, is commonly used to improve the computational efficiency of standard transformers [47]. This pooling technique effectively addresses the issue of representation power (as empirically demonstrated in § 4.7) by expanding the receptive field of the image tokens in the deeper layers. We choose to adopt this approach due to its significant advantages.
@@ 4.2 Implementation Details
To train HiT on ImageNet [16], we used the official DeiT3 codebase [61] and followed a similar setup to their method. Please read the supplementary for more details on the architecture, but in short, HiT uses exactly the same depth and number of layers as ViT. For HiT-B and HiT-S, we trained our models for 600 epochs using the AdamW optimizer [39] with a learning rate of $8\times 10^{-4}$, a weight decay of 0.05, a batch size of 4096, 20 warm-up epochs, a cosine annealing scheduler [38], and ThreeAugment [61] data augmentation. Unlike the DeiT3 training regime, we did not use the binary cross entropy loss or any LayerDrop [21] regularization, but the traditional cross entropy with a smoothing of 0.1 and an attention dropout of 0.2. To fine-tune HiT in the other datasets, we trained our models similarly to the ImageNet’s configuration, but instead we set the batch size to 512, the number of epochs to 300 and the learning rate to $5\times 10^{-5}$ for Standford Dogs and Oxford Pets, $4\times 10^{-4}$ for Standford Cars and FGVC-Aircraft, and $1\times 10^{-4}$ for CUB 2011.
@@ 5 Conclusions
Appendix A Model architecture details
To compare our architecture as fairly as possible with standard methods (i.e., DeiT [60]), we follow the same hyperparameter selection for both the base and small versions as their ViT equivalents. That is, HiT base (HiT-B) follows the same architecture as ViT-B. Similarly, HiT small (HiT-S) follows the same hyperparameters as ViT-S. The only differences are the pooling layers, the initial patch size, and that we have removed the last MLP block as it is not used. Tab. 4 shows the architecture hyperparameter choices compared to Visual Transformers.
Version Layers Feature dimension Heads Pooling layer location Patch dimension Parameter count
HiT-B 12 768 12 $[4,8]$ 8 81.8
ViT-B 12 768 12 - 16 86.9
HiT-S 12 384 6 $[4,8]$ 8 20.8
ViT-S 12 384 6 - 16 22.1
Table 4: HiT base and small hyperparameter configurations
Appendix B Datasets
@@
Appendix C Insertion-Deletion Curves
@@
Appendix D Evaluating HiT without Pooling Layers
We conducted experiments similar to those in the main manuscript to analyze the positive and negative impact of removing pooling layers. As demonstrated in § 4.7 of the paper, pooling layers are essential for improving top-1 accuracy performance. However, their inclusion increases the size of the explanations. First, we explore this phenomenon quantitatively in sections D.1 and § D.2. Later, we will explore the qualitative differences in § D.3.
D.1 Interpretability Trade-off
@@
D.2 Layer-wise Contributions